# Enhancing Code LLM Training with Programmer Attention

### v1

@@ Abstract.
conference: ACM ACM International Conference on the Foundations of Software Engineering; 23-27 June, 2025; Trondheim, Norwaybooktitle: Companion Proceedings of the 33rd ACM Symposium on the Foundations of Software Engineering (FSE ’25), June 23–27, 2025, Trondheim, Norway
@@ 1. Introduction
Meanwhile, LLMs such as CodeT5 have shown substantial gains when fine-tuned on large corpora of code. In human-centered AI for software engineering, recent studies suggest that integrating visual attention patterns (collected via eye tracking) and other developer insights can further improve these models (Zhang et al., 2024; Kou et al., 2024). However, existing data-augmentation methods typically overlook the unique demands of real-world code or the crucial cognitive science underpinnings needed to capture deeper semantic context (Hamza et al., 2024; Chen et al., 2021). As a result, even state-of-the-art LLMs may miss opportunities for further enhancement through programmer attention information (for example, recurring eye-tracking scanpaths or consistent reading orders), particularly when such data are limited. Bridging this gap calls for specialized training pipelines that effectively embed developer attention signals into LLM-based code intelligence.
In this work, we propose HumanLLM, a two-part approach geared primarily toward code summarization, yet we also evaluate its broader impact on completion and translation to test generality. First, we use real eye-tracking data to augment code with developer attention that reflect how programmers read and interpret source files. Second, we embed these signals into a reward-based fine-tuning of CodeT5 (Wang et al., 2021a), aligning the model with genuine programmer fixations. By systematically extracting and reinforcing these human-centric patterns, our method bridges the gap between cognitive insights and machine-driven automation across multiple code tasks.

### v2

@@ Abstract.
Large Language Models, Eye Tracking, Code Comprehension
conference: ACM ACM International Conference on the Foundations of Software Engineering; 23-27 June, 2025; Trondheim, Norwaybooktitle: Companion Proceedings of the 33rd ACM Symposium on the Foundations of Software Engineering (FSE ’25), June 23–27, 2025, Trondheim, Norwaycopyright: rightsretainedjournalyear: 2025conference: 33rd ACM International Conference on the Foundations of Software Engineering; June 23–28, 2025; Trondheim, Norwaybooktitle: 33rd ACM International Conference on the Foundations of Software Engineering (FSE ’25), June 23–28, 2025, Trondheim, Norwaydoi: 10.1145/3696630.3728510isbn: 979-8-4007-1276-0/25/06ccs: Software and its engineering Automatic programmingccs: Information systems Document representation
@@ 1. Introduction
While LLMs like CodeT5 improve when fine-tuned on large code corpora, integrating human signals like eye-tracking attention patterns can enhance them further (Zhang et al., 2024; Kou et al., 2024). However, typical data augmentation often overlooks code’s cognitive demands or specific structures (Hamza et al., 2024; Chen et al., 2021). Consequently, even advanced LLMs may underutilize valuable programmer attention information (like scanpaths or reading orders), especially when such data is sparse. Effectively embedding these developer signals necessitates specialized training pipelines. To address this, we propose HumanLLM, a two-part approach primarily for code summarization (also tested on completion and translation). First, HumanLLM uses real eye-tracking data reflecting programmer reading to augment code. Second, it embeds these signals via reward-based fine-tuning of CodeT5 (Wang et al., 2021a), aligning the model with human fixations. This systematic use of human-centric patterns aims to bridge cognitive insights and automated code intelligence across tasks.